# Optimal Sizing and Material Choice for Additively Manufactured Compact Plate Heat Exchangers

### v1

@@ 4.2 Influence of design effectiveness on optimal heat exchanger
Figure 4: Maximum non-dimensional power density with varying effectiveness (fouling constraint active if filled, inactive if open)
As expected, the non-dimensional power density increases as the design effectiveness decreases for all material cases, indicating that lower effectiveness results in smaller heat exchangers. While this increase in power density is relatively linear across the effectiveness range for low thermal conductivity materials, it becomes more pronounced for higher thermal conductivity materials. This suggests that the dimensions of high thermal conductivity heat exchangers are more sensitive to changes in design effectiveness compared to low thermal conductivity heat exchangers. As a result, the fouling constraint, which is inactive for high thermal conductivity heat exchangers at high effectiveness levels, becomes active at low effectiveness levels due to the reduction in plate spacing (see copper, aluminum, and AlN heat exchangers in Figure 4).
Figure 4 further demonstrates that, while plastic consistently achieves the highest power density across the entire range of design effectiveness values considered, non-dimensional power density decreases as material thermal conductivity increases. Notably, the optimal material choice remains unaffected by changes in design effectiveness.
@@ 5 Impact of the plate thickness on maximal power density
The previous section determined the plate thickness based on AM limits for metal and ceramic heat exchangers. However, these limits can be significantly exceeded with conventional manufacturing (CM) techniques. For instance, the literature reports microchannels made from stainless steel plates as thin as 0.051 mm [44], or even copper plates as thin as 0.003 mm [45], achieved through hydroforming and pressing, respectively. Such thicknesses are far beyond the level of precision currently achievable with AM. This raises the question of whether the optimal material choice changes if AM resolution is improved, or whether current AM methods will continue to act as a bottleneck in cHEX design.
To address this question, optimization studies are conducted for plastic, austenitic steel, and copper with varying minimal plate thicknesses while maintaining the fouling constraint at various design effectiveness levels. The range of plate thicknesses used in this section is set around the material-specific values from the previous section. For plastic, the material-specific thickness was determined based on structural concerns rather than the limitations of AM. Nonetheless, the same procedure is applied to the plastic heat exchanger to examine the effect of plate thickness on its power density. The results of these optimizations are depicted in Figure 5. Each point in this figure corresponds to an optimization run.
@@
Despite this, due to copper’s higher thermal conductivity, it still results in a larger heat exchanger compared to austenitic steel and plastic, as shown in Figure 5. When the same plate thickness of 0.1 mm or 0.2 mm is used in the optimizations for all materials, copper consistently exhibits the lowest power density across all effectiveness values. However, the austenitic steel heat exchanger achieves the highest power density for all effectiveness values except 0.94. This indicates that advancements in austenitic steel thin-wall production through AM could make austenitic steel competitive with, or even superior to, plastic heat exchangers in terms of compactness.
@@
Figure 5: Non-dimensional power density variation with plate thickness $t_{d}$ (fouling constraint active if filled, inactive if open)
@@ 6 Discussion
This study demonstrates that materials with lower thermal conductivity, such as plastics and ceramics, exhibit greater potential for maximizing compactness in cHEX design compared to metals like copper or aluminum. This is primarily attributed to the dominance of axial heat conduction in highly thermally conductive materials, which diminishes effectiveness in cHEXs. Consequently, for a given effectiveness, heat exchangers made from materials with higher thermal conductivity need to be larger than those made from lower thermal conductivity materials. However, the final compactness of a cHEX is influenced not only by material properties but also by manufacturing and fouling constraints, such as the minimum achievable plate thickness and required plate spacing.
@@
In optimizations conducted without manufacturing or fouling constraints at $\varepsilon_{d}$ = 0.79, plastics outperform other materials in terms of power density, achieving an improvement factor of 1785.91 compared to the initial heat exchanger. Although fixing the plate thickness at 0.5 mm for all materials reduces this difference, plastic still achieves the highest power density among the materials, with an improvement factor of nearly 4.0 compared to the additive manufacturable initial steel heat exchanger. However, when the fouling constraint is introduced at the same fixed plate thickness, austenitic steel emerges as the material with the highest power density. These findings align with the existing literature. Stief et al. [11] suggest that the high thermal conductivity of cHEX plates negatively impacts thermal performance, while very low thermal conductivity can act as insulation, also reducing thermal performance.
@@
When the fixed plate thickness of 0.5 mm is replaced with material-specific plate thicknesses to represent the varying resolution capabilities in AM, the plastic heat exchanger once again achieves the highest power density. This is primarily because AM thin-wall production has not yet reached the desired precision for metal and ceramic materials. With material-specific plate thicknesses, the power density of the plastic heat exchanger is 5.06 times higher than that of the optimal copper heat exchanger. Similarly, the power density of austenitic steel is 4.42 times higher, Al2O3 is 4.34 times higher, AlN is 2.91 times higher, and aluminum is 2.12 times higher than copper. This means that even the aluminum heat exchanger achieves nearly half the volume of the copper heat exchanger in this optimization scenario. This outcome is driven not only by the increased volume associated with thicker plates but also by the effect of thicker plates enhancing axial heat conduction. This increased conduction necessitates longer heat exchanger lengths and wider plate spacings to achieve the desired design effectiveness.
@@
The observation that the copper heat exchanger has the lowest power density compared to other materials holds true for all $\varepsilon_{d}$ in the range of 0.55 to 0.94 when using material-specific plate thicknesses, as shown in Figure 4. Meanwhile, plastic consistently outperforms all other materials across this range. However, as expected, the power density decreases with rising effectiveness for all materials, indicating the increase in heat exchanger dimensions with effectiveness. The increase in plate spacing with effectiveness can be observed by examining the activation of the fouling constraint in the same figure. At high effectiveness values, particularly for plates with high thermal conductivity, the fouling constraint becomes inactive. This occurs because the maximum power density is achieved with a plate spacing configuration exceeding the 0.8 mm minimum requirement.
Another critical aspect investigated in this study is the effect of plate thickness on the power density of cHEXs, as given in Section 5. The optimizations in this section evaluate how advancements in AM resolution could enhance power density or influence the optimal material selection for cHEX designs.
@@
Figure 5 demonstrates that the limitations of AM in producing thin features constrain the maximum power density. Reducing plate thickness increases the maximum power density across all examined materials, with the effect being most pronounced for copper compared to plastic and austenitic steel. These results highlight the potential for advancements in thin-wall manufacturing to significantly improve the compactness of copper heat exchangers. Despite this potential, copper’s high thermal conductivity still results in larger heat exchangers compared to austenitic steel and plastic. Even with plate thicknesses of 0.1 mm or 0.2 mm, copper consistently achieves the lowest power density across all effectiveness values.
Interestingly, when the plate thickness is set to 0.1 mm or 0.2 mm, austenitic steel achieves the highest power density for all effectiveness values except 0.94. At an effectiveness of 0.6, its power density surpasses that of plastic by factors of 1.06 and 1.12 for plate thicknesses of 0.1 mm and 0.2 mm, respectively. However, at an effectiveness of 0.92, the power densities of austenitic steel and plastic are nearly identical for both plate thickness values. Finally, at an effectiveness of 0.94, the power density of austenitic steel falls below that of plastic by factors of 0.98 and 0.96 for plate thicknesses of 0.1 mm and 0.2 mm, respectively. While these findings highlight the potential of austenitic steel, these plate thickness values exceed the current capabilities of AM technology for producing stable, leak-proof austenitic steel heat exchangers.
@@
This study focused on a counterflow PFPHE. However, the results obtained for maximum power density without plate thickness and fouling constraints were compared with those of a crossflow PFPHE. The analysis of crossflow PFPHE utilized the correlations presented by Buckinx et al. [28]. For the crossflow heat exchangers, the same design effectiveness value of 0.79 was used, and the maximum power density, along with the corresponding optimal length and plate spacing for the same six materials, was calculated. The results showed that the inverse relationship between maximum power density and thermal conductivity observed in the counterflow case was maintained. However, for each material, crossflow heat exchangers exhibited approximately nine times lower maximum power density compared to counterflow heat exchangers.
@@
Finally, it is crucial to note that although plastic yields the smallest volume heat exchanger for most of the optimization scenarios, its susceptibility to high temperatures limits its practical applicability.
@@ 7 Conclusions
This study investigates the influence of thermal material properties and AM constraints on the sizing of compact heat exchangers (cHEXs). The findings reveal that, for a given effectiveness, a typical plastic, austenitic steel, copper, aluminum, aluminum nitride (AlN), and aluminum oxide (Al2O3) enable varying levels of compactness. Among these, plastic allows for the design of significantly more compact heat exchangers compared to the other materials when manufacturing and fouling constraints are not considered. Even with a fixed plate thickness constraint of 0.5 mm for all materials, plastic retains the highest power density, achieving 12.2 times greater power density than copper.
However, the absence of fouling constraints results in impractically small plate spacing values in some configurations, such as 0.02 mm for plastic in unconstrained optimization or 0.14 mm with only a plate thickness constraint. These designs are highly prone to fouling. Introducing a fouling constraint shifts the optimal material to austenitic steel when the same manufacturable plate thickness is applied across all materials.
Copper, despite its high thermal conductivity, consistently exhibits the lowest power density in all optimization strategies due to significant performance degradation caused by axial conduction in the wall. This results in less compact designs. High thermal conductivity materials are further limited by the relatively lower resolution of AM techniques for producing thin walls. Thicker plate requirements worsen axial conduction effects and cause further increases in heat exchanger dimensions when maintaining constant effectiveness.
Using material-specific plate thicknesses in the optimizations accounts for AM’s varying capabilities for thin-wall production in different materials. These optimizations confirm that plastic remains the most promising material, benefiting from the relatively higher resolution of AM plastics. This trend is consistent across all effectiveness values considered, with plastic maintaining the highest power density.
@@
This study assumes constant material properties for the fluid and solid, and does not consider the mechanical strength of plates or allowable operational temperatures. For high-temperature applications, the results suggest that austenitic steel may be a favorable choice. Additionally, for simplicity, only unfinned parallel flat plate heat exchangers are investigated. Future research should address these factors to draw more general conclusions on the most suitable materials for AM cHEXs.
@@ 8 Acknowledgment
This work received funding from VLAIO through the SIM/Flux50 cSBO IAMHEX project and from a KU Leuven starting grant with reference STG/21/016.

### v2

@@ 4.2 Influence of design effectiveness on optimal heat exchanger
Figure 3: Maximum nondimensional power density with varying effectiveness (fouling constraint active if filled, inactive if open)
As expected, the nondimensional power density increases as the design effectiveness decreases for all material cases, indicating that lower effectiveness results in smaller heat exchangers. While this increase in power density is relatively linear across the effectiveness range for low thermal conductivity materials, it becomes more pronounced for higher thermal conductivity materials. This suggests that the dimensions of high thermal conductivity heat exchangers are more sensitive to changes in design effectiveness compared to low thermal conductivity heat exchangers. As a result, the fouling constraint, which is inactive for high thermal conductivity heat exchangers at high effectiveness levels, becomes active at low effectiveness levels due to the reduction in plate spacing (see copper, aluminum, and AlN heat exchangers in Figure 3).
Figure 3 further demonstrates that, while plastic consistently achieves the highest power density across the entire range of design effectiveness values considered, nondimensional power density decreases as material thermal conductivity increases. Notably, the optimal material choice remains unaffected by changes in design effectiveness.
@@ 5 Impact of the plate thickness on maximal power density
The previous section determined the plate thickness based on AM limits for metal and ceramic heat exchangers. However, these limits can be significantly exceeded with conventional manufacturing (CM) techniques. For instance, the literature reports microchannels made from stainless steel plates as thin as 0.051 mm Mahabunphachai2008 , or even copper plates as thin as 0.003 mm Joo2004 , achieved through hydroforming and pressing, respectively. Such thicknesses are far beyond the level of precision currently achievable with AM. This raises the question whether the optimal material choice changes if AM resolution is improved, or whether current AM methods will continue to act as a bottleneck in cHEX design.
To address this question, optimization studies are conducted for plastic, austenitic steel, and copper for varying minimal plate thicknesses and design effectiveness levels, while maintaining the fouling constraint fixed. The range of plate thicknesses used in this section is set around the material-specific values from the previous section. For plastic, the material-specific thickness was determined based on structural concerns rather than the limitations of AM. Nonetheless, the same procedure is applied to the plastic heat exchanger to examine the effect of plate thickness on its power density. The results of these optimizations are depicted in Figure 4. Each point in this figure corresponds to an optimization run.
@@
Despite copper’s higher thermal conductivity, it still results in a larger heat exchanger compared to austenitic steel and plastic, as shown in Figure 4. When the same plate thickness of $t^{\ast}=0.63$ ($t=0.1\,\text{mm}$) or $t^{\ast}=1.25$ ($t=0.2\,\text{mm}$) is used in the optimizations for all materials, copper consistently yields the lowest power density across all effectiveness values. In contrast, the austenitic steel heat exchanger achieves the highest power density at effectiveness values of 0.6 and 0.7, while plastic performs better at higher effectiveness levels. This suggests that advances in thin-wall manufacturing of austenitic steel via AM could make it competitive with, or even superior to, plastic in terms of compactness under certain operating conditions.
@@
Figure 4: Nondimensional power density variation with nondimensional plate thickness $t^{\ast}$ (fouling constraint active if filled, inactive if open). The vertical dashed line indicates the material-specific plate thickness used in Section 4.
@@ 6 Discussion
This study demonstrates that materials with lower thermal conductivity, such as plastics and ceramics, exhibit greater potential for maximizing compactness in cHEX design compared to metals like copper or aluminum. This is primarily attributed to the dominance of axial heat conduction in highly thermally conductive materials, which diminishes effectiveness in cHEXs. Consequently, for a given effectiveness, heat exchangers made from materials with higher thermal conductivity need to be larger than those made from lower thermal conductivity materials. However, the final compactness of a cHEX is influenced not only by material properties but also by manufacturing and fouling constraints, such as the minimum achievable plate thickness and the required minimum plate spacing.
@@
In optimizations conducted without manufacturing or fouling constraints at $\varepsilon_{d}$ = 0.79, plastic achieves the highest power density, with an improvement factor of 962 relative to the initial heat exchanger. When a uniform plate thickness of 0.5 mm is applied to all materials, plastic continues to outperform others, reaching an improvement factor of 4.36 compared to the additive manufacturable reference plastic design. However, when the fouling constraint is introduced at the same fixed plate thickness, the optimal plastic design exhibits a lower power density than the reference plastic design. Despite this reduction, plastic still performs better than the other materials, followed closely by austenitic steel. In this optimization scenario, the plastic achieves 1.02 times the performance of austenitic steel.
@@
When material-specific plate thicknesses are applied to reflect the resolution limits of different AM processes, the performance advantage of plastic becomes evident once more, with an improvement factor of 1.52 relative to the additive manufacturable reference plastic design. This is primarily because AM thin-wall production has not yet reached the desired precision for metal and ceramic materials. With material-specific plate thicknesses, the power density of the plastic heat exchanger is 7.5 times higher than that of the optimal copper heat exchanger. Similarly, the power density of austenitic steel is 6.06 times higher, Al2O3 is 5.87 times higher, AlN is 2.81 times higher, and aluminum is 2.06 times higher than copper. This means that even the aluminum heat exchanger achieves nearly half the volume of the copper heat exchanger in this optimization scenario. This outcome is driven not only by the increased volume associated with thicker plates but also by the effect of thicker plates enhancing axial heat conduction. This increased conduction necessitates longer heat exchanger lengths and wider plate spacings to achieve the desired design effectiveness.
@@
The observation that the copper heat exchanger has the lowest power density compared to other materials holds true for all $\varepsilon_{d}$ in the range of 0.55 to 0.94 when using material-specific plate thicknesses, as shown in Figure 3. Meanwhile, plastic consistently outperforms all other materials across this range in terms of power density, though other criteria like operational temperatures might also influence the material selection. However, as expected, the power density decreases with rising effectiveness for all materials, indicating the increase in heat exchanger dimensions with effectiveness. The increase in plate spacing with effectiveness can be observed by examining the activation of the fouling constraint in the same figure. At high effectiveness values, particularly for plates with high thermal conductivity, the fouling constraint becomes inactive. This occurs because the maximum power density is achieved with a plate spacing configuration exceeding the 0.8 mm minimum requirement.
These trends can also be understood in light of the asymptotic high-NTU effectiveness limit. For counterflow arrangements this limit is given by
$\varepsilon_{\mathrm{NTU}\rightarrow\infty}=\frac{\textup{M}+1}{2\textup{M}+1},$
with M defined in Eq. 12. This expression shows that, independent of channel length, the attainable effectiveness may be capped below the design target if M is sufficiently large. The controlling manufacturing property in this regard is the product $k_{w}t_{\text{min}}$, i.e., the material thermal conductivity multiplied by the minimum manufacturable wall thickness. For high-conductivity materials, AM processes generally require larger wall thicknesses, and this combination restricts the effectiveness achievable in compact designs. While such limitations could be alleviated by allowing for higher pressure drops or larger plate spacings, the latter comes at the expense of compactness and illustrates an inherent trade-off that makes very high conductivity materials less suitable for small flow passages in additively manufactured heat exchangers. These trade-offs further emphasize that plate thickness plays a crucial role in determining the maximum achievable power density, as pointed out in Section 5.
@@
The optimizations in Section 5 evaluate how advancements in AM resolution could enhance power density or influence the optimal material selection for cHEX designs. Figure 4 demonstrates that the limitations of AM in producing thin features constrain the maximum power density. Reducing plate thickness increases the maximum power density across all examined materials, with the effect being most pronounced for copper compared to plastic and austenitic steel. These results highlight the potential for advancements in thin-wall manufacturing to significantly improve the compactness of copper heat exchangers. Despite this potential, copper’s high thermal conductivity still results in larger heat exchangers compared to austenitic steel and plastic. Even with plate thicknesses of $t^{\ast}=0.63$ ($t=0.1\,\text{mm}$) or $t^{\ast}=1.25$ ($t=0.2\,\text{mm}$), copper consistently achieves the lowest power density across all effectiveness values.
Interestingly, when the plate thickness is set to 0.1 mm or 0.2 mm, austenitic steel outperforms other materials in power density for effectiveness values of 0.6 and 0.7. These results align with previous findings in the literature. Stief et al. High thermal conductivity in cHEX plates can impair thermal performance due to increased axial conduction, while excessively low conductivity may lead to insulation-like behavior, also diminishing performance Stief1999 . At an effectiveness of 0.6, the power density of austenitic steel exceeds that of plastic by factors of 1.02 and 1.04 for plate thicknesses of 0.1 mm and 0.2 mm, respectively. However, as the target effectiveness increases to 0.79, the power densities of plastic and austenitic steel converge and become nearly identical for both thickness values. At higher effectiveness levels, plastic consistently outperforms austenitic steel. While these findings highlight the potential of austenitic steel in specific low-effectiveness scenarios, the thin plate values required for such performance still exceed current AM capabilities for producing structurally robust, leak-tight steel heat exchangers.
@@
This study focuses on a counterflow PFPHE. However, the results obtained for maximum power density without plate thickness and fouling constraints are also compared with those of a crossflow PFPHE. The analysis of crossflow PFPHE utilized the correlations presented by Buckinx et al. Buckinx2013 . For the crossflow heat exchangers, the same design effectiveness value of 0.79 was used, and the maximum power density, along with the corresponding optimal length and plate spacing for the same six materials, was calculated. The results showed that the inverse relationship between maximum power density and thermal conductivity observed in the counterflow case was maintained. However, for each material, crossflow heat exchangers exhibited approximately nine times lower maximum power density compared to counterflow heat exchangers.
@@
Finally, it is crucial to note that although plastic yields the smallest volume heat exchanger for most of the optimization scenarios, its susceptibility to high temperatures might limit practical applicability.
@@ 7 Conclusions
This study investigates the influence of additive manufacturing (AM) and fouling constraints on the sizing of compact heat exchangers (cHEXs) for different materials. The analysis is based on a nondimensional thermal model that captures the coupled effects of fluid flow, heat transfer, and axial conduction, and its accuracy is validated through CFD simulations, showing close agreement with analytical predictions. This model is applied in numerical optimization to evaluate power density under additive manufacturing and fouling constraints for 6 different materials.
The findings reveal that, for a given effectiveness, materials such as a representative low-conductivity plastic, austenitic steel, copper, aluminum, aluminum nitride (AlN), and aluminum oxide (Al2O3) enable varying levels of compactness. Among these, plastic consistently enables more compact heat exchanger designs in the majority of optimization scenarios. The greatest performance advantage occurs when manufacturing and fouling constraints are not applied, yielding an improvement factor (IF) of 962 relative to the baseline design. Even when all materials are subject to a fixed plate thickness constraint of 0.5 mm, the plastic retains the highest power density, achieving 22 times greater performance than the optimal copper heat exchanger.
[…]
Copper, despite its high thermal conductivity, consistently exhibits the lowest power density in all optimization strategies due to significant performance degradation caused by axial conduction in the wall. This results in less compact designs. High thermal conductivity materials are further limited by the relatively lower resolution of AM techniques for producing thin walls. Thicker plate requirements worsen axial conduction effects and cause further increases in heat exchanger dimensions when maintaining constant effectiveness. Using material-specific plate thicknesses in the optimizations accounts for AM’s varying capabilities for thin-wall production in different materials. These optimizations confirm that plastic remains the most promising material, benefiting from the relatively higher resolution of AM plastics. This trend is consistent across all effectiveness values considered, with plastic maintaining the highest power density.
@@
This study assumes constant material properties for the fluid and solid, and does not consider the mechanical strength of the plates or allowable operating temperatures. For high-temperature applications, the results suggest that austenitic steel may be a favorable choice. Additionally, for simplicity, only unfinned counterflow parallel flat plate heat exchangers are investigated. Future research should address these factors to draw more general conclusions on the most suitable materials for AM cHEXs, and may also extend the present analysis to other flow arrangements, such as cross-flow configurations.
@@ 8 Acknowledgment
This work received funding from VLAIO under project number HBC.2021.0801/IAMHEX and from a KU Leuven starting grant with reference STG/21/016.